Two small conducting and identical spheres A and B have charges -25 nC and +15 nC, respectively. They are separated by a distance of 0.02 m. (a) What is the magnitude of the electric force between the two spheres? Is this force attractive or repulsive? (b) The spheres are then allowed to touch each other and then separated. What is the magnitude of the force between the two spheres? Is this a repulsive force or an attractive force?


1
Expert's answer
2022-02-14T17:01:07-0500

Part a)


"F=k\frac{q_1q_2}{r_2}"


"=\frac{9×10^9×25×15×(10^{-9})^2}{(0.02)^2}"


"=8.438\u00d710^{-3}N"


They have opposite charges

"\implies" The force is attractive



b)

Total charge "=-25+15=-10nC"

Each attains "=\frac{-10}{2}"

"=-5nC"


If they are separated at initial distance of "0.02m"


"F=\frac{9×10^9×5×5×(10^{-9})^2}{(0.02)^2}"


"=5.625\u00d710^{-4}N"


They have same charge.

Therefore the force is repulsive